OpenStudy (anonymous):

solve by substitution method 0.08x-0.01y=4.4 0.26x+0.14y=21.2

OpenStudy (anonymous):

(1) 0.08x-0.01y=4.4 (2) 0.26x+0.14y=21.2 From (1): 0.08x=4.4+0.01y x = (4.4+0.01y)/0.08 (1.1) Substituting into (2): 0.26x+0.14y=21.2 0.26[ (4.4+0.01y)/0.08 ] + 0.14y = 21.2 ( 1.144 + 0.0026y ) / 0.08 + 0.14y = 21.2 14.3 + 0.0325y + .14y = 21.2 0.1725y = 6.9 y = (6.9/0.1725) y = 40 ans. Plug into (1.1): x = (4.4+0.01(40))/0.08 x = 60 phew

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you could have just multiplied both eqns by 100 , that would have eliminated the messy decimals